A density of bottle contains 104.o gm of a liquid at 34(celsius temperate )on heating it at 100(c t) the mass of liquid expelled is 4.5 gm . then find the coefficient of apparent expansion of liquid?

Grade:11

2 Answers

sivakrishna
10 years ago
Let the density of the liquid in the bottle be d. Given that, m1=104gm. t1=34celsius m2=m1+4.5 =104+4.5 =108.5 t2=100celsius we know that the coefficient of apparent expansion=(V2-V1)/V1.(t2-t1) =(m2/d-m1/d)/m1/d(t2-t1) =(108.5/d-104/d)/104/d(66) =(4.5/d)/104*66/d =4.5/(104*66) =4.5/6864 =6.55*10^-4
Shaik Irfan Kareem Ali
10 years ago
here the density of liquid is constant we know ??V=V1????t ??V/V=????t we also know d=m/V since d is constant , ??V/V=??m/m ??m=4.5gm m=104gm ??t=66°C ??=(??V/V)/??t ??=(??m/m)/??t =(4.5/104)/66 =6.5559*10?4 ??=0.0006559 /°C approx. ?? = 0.0007 /°C